OpenStudy (anonymous):

The legs of a triangle measure 2(sqrt 3) and 5. What is the length of hypotenuse? Help

OpenStudy (anonymous):

sqrt(12+25)

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[2\sqrt{3}\]

OpenStudy (anonymous):

sqrt37

OpenStudy (anonymous):

(2√3)^2=12

OpenStudy (anonymous):

\[\sqrt{(2\sqrt{3})^2+5^2=4\times 3 + 25} = \sqrt{12+25}=\sqrt{37}\]